Jorge's brother is 4 less than 3 times his age. The sum of their ages is 64. How old is Jorge's brother?

Answer by stanbon(75887)

Jorge's brother is 4 less than 3 times his age. The sum of their ages is 64. How old is Jorge's brother?

----------------
Equations:
b = 3j - 4
b + j = 64
--------------------
Substitute for "b" and solve for "j"::
3j-4 + j = 64
4j = 68
j = 17 yrs (Jorge's age now)
b = 3j-4 = 47 yrs (His brother's age now)
==============
Cheers,
Stan H.
===============